7.30 L container holds a mixture of two gases at 43 °C.

43 °C. The partial pressures of gas A and gas B, respectively, are 0.272 atm

0.272 atm and 0.859 atm.

0.859 atm. If 0.210 mol

0.210 mol of a third gas is added with no change in volume or temperature, what will the total pressure become?


1
Expert's answer
2021-11-11T12:17:02-0500

PV=nRT

R=0.0821(L*atm)/(K*mol)

T=51+273=324K

V=7.30L

 

Total Pressure equals the sum of the partial pressures of the gasses in the container so initial pressure is sum of the pressure of gas A and gas B.

 

P1=0.308atm+0.520atm=0.828atmSolve for initial moles of gas

n1= (0.828*8.1)/(0.0821*324)mol Add 0.120mol of gas to n1 to get n2 and use nto solve for P2 or the total pressure once the third gas has been added